# MixNet: Towards Effective and Efficient UHD Low-Light Image Enhancement

### v1

@@ 4.3 Ablation Study
Table 3: Comparison of quantitative results on O-Haze dataset. Best and second best values are indicated with bold text and underlined text respectively.
Task Methods PSNR SSIM
non-UHD DCP He et al. (2010) 14.54 0.333
BCCR Meng et al. (2013) 15.20 0.361
CAP Zhu et al. (2015) 15.35 0.342
NLD Berman et al. (2016) 14.76 0.396
DehazeNet Cai et al. (2016) 12.39 0.359
AodNet Li et al. (2017) 15.87 0.581
GridNet Liu et al. (2019) 17.44 0.603
PMSNet Chen et al. (2019) 16.99 0.597
MSBDN Dong et al. (2020) 17.46 0.635
UHD MGBL Zheng et al. (2021a) 18.72 0.614
NSEN Yu et al. (2023) 19.23 0.672
MixNet 21.46 0.650
4.4 Extension Experiments
Although we primarily focus on the UHD-LLIE task in this paper, it is important to note that we do not introduce specific prior knowledge related to low-light conditions. Therefore, our method can be considered as a general approach that performs well not only in UHD-LLIE task but also in various UHD image restoration task, such dehazing.
4.4.1 Datasets
We conduct dehazing experiments on the O-HAZE dataset which has 45 paired hazy and haze-free images. Following the prior work Yu et al. (2023), we randomly choose 35 images for training and the others for testing.
4.4.2 Results
We compare the proposed MixNet with several image dehazing approaches Berman et al. (2016); He et al. (2010); Meng et al. (2013); Zhu et al. (2015); Li et al. (2017); Dong et al. (2020); Liu et al. (2019); Cai et al. (2016); Chen et al. (2019); Zheng et al. (2021a); Yu et al. (2023). Results are reported in Tab. 3. MixNet achieves a substantial gain of over 2 dB compared to NSEN in PSNR on the UHD image dehazing task.
Table 4: Ablation study of proposed blocks. In order to ensure consistency in parameter magnitude, we gradually replace the proposed module with Residual Block of comparable parameters.
# GFML LFML FFL PSNR SSIM Parameter
@@ 4.4.2 Results
Figure 7: Visualizations of the feature maps before the UpSampler. The image used here is the low-light image shown in Fig. 6. Our method reconstructs feature maps with sharper textures.
5 Conclusion
In this paper, we proposed a simple yet efficient framework called MixNet to solve the efficient UHD-LLIE problem. MixNet has a series of FMBs which mainly contain a GFML, LFML, and FFL. GFML explored long-range dependency modeling upon the simple dimensional transformation operation. By permuting the feature maps, GFML associates features from different views. LFML and FFL are designed to capture local features and transform features into a compact representation, respectively. Extensive experiments on both synthetic and real-world datasets showed that the proposed MixNet is more efficient than state-of-the-art methods while achieving competitive performance.

### v2

@@ Introduction
Table 1: Comparison of quantitative results on UHD-LOL4K dataset. Best and second best values are indicated with bold text and underlined text respectively.
Experiments
Implementation Details and Datasets
Implementation Details.
We conduct experiments in PyTorch on six NVIDIA GeForce RTX 3090 GPUs. To optimize the network, we employ the Adam optimizer with a initial learning rate $2\times 10^{-4}$ and a cosine annealing strategy is used for the decay of the learning rate. We randomly crop the full-resolution 4K image to a resolution of $512\times 512$ as the input and the batch size is set ot 24. Except for the UHD underwater image enhancement task, where the dataset is relatively small, we only perform 100K iterations. For all other tasks, we perform 300K iterations. To augment the training data, random horizontal and vertical flips are applied to the input images. The number of FMB and feature channels is set to 8 and 48, respectively.
Datasets.
To verify the effectiveness of MixNet, we evaluate it on the UHD-LOL dataset (Wang et al. 2023), UHD-Blur dataset (Wang et al. 2024), UHD-UW dataset (Wei, Zheng, and Jia 2022) and UHDM dataset (Yu et al. 2022a). UHD-LOL consists of two subsets, namely UHD-LOL4K and UHD-LOL8K, which contain UHD images with resolutions of 4K and 8K in low-light conditions, respectively. In this study, we focus on the UHD-LOL4K subset to validate the effectiveness of WowNet. The UHD-LOL4K subset comprises a total of 8,099 image pairs, with 5,999 pairs allocated for training purposes and 2,100 pairs designated for testing. UHD-Blur dataset is re-collected from the datasets of Deng et al. (Deng et al. 2021) by Wang et al. (Wang et al. 2024) and it has 1,964 blur-clear image pairs for training and 300 pairs for testing. UHD-UW dataset is our organization of the dataset proposed in UHD-SFNet (Wei, Zheng, and Jia 2022), where we remove the images with resolutions lower than $512\times 512$, and sequentially select 80 images as our training set, with the remaining 13 images as our test set. UHDM dataset is a real-world 4K resolution demoir´eing dataset, consisting of 5,000 image pairs in total, with 4500 images used for training and 500 images used for testing.
Methods Type Venue PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
Shallow-uwnet non-UHD AAAI’21 23.36 0.798 0.22M
USUIR AAAI’22 24.13 0.834 3.36M
PUIE-Net ECCV’22 25.77 0.859 1.40M
URank AAAI’23 26.39 0.911 3.15M
FA+Net BMVC’23 27.60 0.916 9K
UHD UHD ICCV’21 26.89 0.907 34.5M
UHD-SFNet ACCV’22 26.31 0.914 37.31M
Ours - 28.09 0.932 7.77M
Table 2: Comparison of quantitative results on UHD-UW dataset. Best and second best values are indicated with bold text and underlined text respectively.
Comparisons with the State-of-the-art Methods
Evaluation.
Two well-known metrics, Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) (Wang et al. 2004), are employed for quantitative comparisons. Higher values of these metrics indicate superior performance of the methods. PSNR and SSIM are calculated along the RGB channels. We also report the trainable parameters (Param) and running time (RT) which you can find in Tab. 7. For methods that cannot perform full-resolution inference, we adopt the previous patch-based or downsample-based approaches to handle them.
Figure 6: Visual quality comparisons with state-of-the-art methods on UHD-Blur dataset. Please zoom in for details.
Methods Type Venue PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
MIMO-Unet++ non-UHD ICCV’21 25.03 0.752 16.1M
Restormer CVPR’22 25.21 0.752 26.1M
Uformer CVPR’22 25.27 0.752 20.6M
Stripformer ECCV’22 25.05 0.750 19.7M
FFTformer CVPR’23 25.41 0.757 16.6M
UHDformer UHD AAAI’24 28.82 0.844 0.34M
Ours - 29.43 0.855 5.22M
Table 3: Comparison of quantitative results on UHD-Blur dataset. Best and second best values are indicated with bold text and underlined text respectively.
Low-light UHD Image Enhancement Results.
For low-light UHD image enhancement task, we compare our MixNet with approaches Z_DCE++ (Li, Guo, and Loy 2021), RUAS (Liu et al. 2021), ELGAN (Jiang et al. 2021), Uformer (Wang et al. 2022), Restormer (Zamir et al. 2022), NSEN (Yu et al. 2023), UHDFour (Li et al. 2023a), LLFormer (Wang et al. 2023) and UHDformer (Wang et al. 2024). As shown in Tab. 1, our method outperforms the current state-of-the-art method, LLFormer, by 1.89dB in PSNR, which is undoubtedly a significant improvement. And you can find the visualization results in Fig. 4. From the color histograms, it can be seen that the images restored using our method have the color distribution closest to the ground truth.
Underwater UHD Image Enhancement Results.
We evaluate underwater UHD image enhancement results on UHD-UW dataset and compare our method with several methods (Naik, Swarnakar, and Mittal 2021; Fu et al. 2022a, b; Guo et al. 2023; Jiang et al. 2023; Zheng et al. 2021a; Wei, Zheng, and Jia 2022). The results are shown in Tab. 5, and compared with recent state-of-the-art FA+Net (Jiang et al. 2023), MixNet achieves a 0.49dB performance improvement in PSNR. You can see the comparison of the visualization results in Fig. 5, which shows that our method achieves better color correction.
UHD Image Deblurring Results.
In the UHD image deblurring task, we evaluate our proposed MixNet against existing deblurring approaches, including MIMO-Unet++ (Cho et al. 2021), Restormer (Zamir et al. 2022), Uformer (Wang et al. 2022), Stripformer (Tsai et al. 2022), FFTformer (Kong et al. 2023) and UHDformer (Wang et al. 2024). As shown in Tab. 3, our method achieved a 0.61 dB improvement in PSNR compared to the current state-of-the-art method, UHDformer. Additionally, you can find the visual comparison results in Fig. 6.
UHD Image Demoiréing Results.
To validate the effectiveness of our method on the task of moiré pattern removal, we compare it with many methods, including MDDM (Cheng, Fu, and Yang 2019), MopNet (He et al. 2019), WDNet (Liu et al. 2020), MBCNN (Zheng et al. 2020), FHDe2Net (He et al. 2020) and ESDNet (Yu et al. 2022b). As shown in Tab. 4, our method achieves a 0.23 dB improvement in PSNR compared to the ESDNet, which is also specifically designed for UHD image demoiréing. You can find the comparison of visualization results in Fig. 7.
Methods Type Venue PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
MDDM non-UHD ICCVW’19 20.09 0.744 7.63M
MopNet ICCV’19 19.48 0.757 58.5M
WDNet ECCV’20 20.36 0.649 3.36M
MBCNN CVPR’20 21.41 0.793 14.19M
FHDe2Net ECCV’20 20.33 0.749 13.5M
ESDNet UHD ECCV’22 22.11 0.795 5.93M
Ours - 22.34 0.799 7.77M
Table 4: Comparison of quantitative results on UHDM dataset. Best and second best values are indicated with bold text and underlined text respectively.
Figure 7: Visual quality comparisons with state-of-the-art methods on UHDM dataset. Please zoom in for details.
Ablation Study
We further conduct extensive ablation studies to better understand and evaluate each component in the proposed MixNet. For fair comparisons with the designed baselines, we implement all experiments with the same setting on UHD-LOL4K dataset.
# GFML LFML FFL PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
@@
Table 5: Ablation study of proposed blocks. In order to ensure consistency in parameter magnitude, we gradually replace the proposed module with ResBlock.
Stage1 Stage2 Stage3
PSNR $\uparrow$ 37.94 38.57 39.22
SSIM $\uparrow$ 0.989 0.990 0.992
Table 6: Ablation study of the number of permute stages. We gradually remove some stages to verify their necessity.
Effectiveness of Proposed Modules.
To demonstrate the gains of proposed modules, we gradually replace the proposed modules with the Residual Block (ResBlock) (He et al. 2016) of comparable parameters, and the experimental results in Tab. 5. Based on all the results in the Tab. 5, all the key designs contribute to the best performance of the full model. Without the GFML (#a, #c, #d, #f), the model shows the most severe performance degradation. This aligns with our hypothesis, as the excellent long-range dependency modeling capability brought by GMFL is one of the key factors contributing to the outstanding performance of MixNet. And the result suggests the importance of long-range dependency modeling among features. From the results of #g and #e, we find that the impact of FFL on performance is higher than that of GFML. We believe that this might be the ability of the FFL to learn compact feature representations, which the ResBlock lacks. And any CNN is capable of learning local features. From the results of #b, #c and #d, we observe that when using the block we proposed alone, GFML and LFML have a higher impact on the results than FFL. This may be because when the feature representation is not effective enough (not from GFML and LFML), learning its compact representation has little effect.
Effectiveness of the Number of Permute Stages.
GFML is the core part of our method. As can be clearly seen in Fig. 3, its main process can be divided into three steps. Here we verify the necessity of these three stages. In Tab 6, we can see that the network’s performance drops sharply when some stages are removed. We believe this is because when part of the permute operation is removed, the global modeling capability of GFML is affected, and the remaining permute operation is not enough to complete the long-range dependency modeling of the features.
UHD LLFormer UHDFour UHDformer Ours
Param $\downarrow$ 34.5M 24.52M 17.54M 0.34M 7.70M
RT $\downarrow$ 1.58s 7.12s 1.67s 5.85s 0.53s
Table 7: Trainable parameters and running time of some UHD image restoration methods. We record the time cost required to perform inference on $3840\times 2160$ size images using a single NVIDIA GeForce RTX 3060 GPU.
Figure 8: Examples of unsatisfactory restorations. Please zoom in for details.
Conclusion
In this paper, we propose a simple yet effective framework called MixNet to solve the efficient UHD image restoration problem. MixNet has a series of FMBs which mainly contain a GFML, LFML, and FFL. GFML explores long-range dependency modeling upon the simple dimensional transformation operation. By permuting the feature maps, GFML associates features from different views. LFML and FFL are designed to capture local features and transform features into a compact representation, respectively. Extensive experiments on both synthetic and real-world datasets showed that the proposed MixNet is more efficient than state-of-the-art methods while achieving competitive performance.
Limitations.
Although our proposed MixNet achieves satisfactory performance across multiple tasks, its efficacy remains limited when dealing with real-world datasets containing complex degradation patterns. For example, in the task of moiré pattern removal, as shown in Fig. 8, the complex moire patterns have not been effectively removed, and the visual results still have a certain gap from the ground truth.